OpenStudy (anonymous):

help?!? what is the degree of the polynomial? 9b^4 c^3 - 6b^3 c4

OpenStudy (whpalmer4):

the degree of the polynomial is the highest degree of any of the terms. The degree of a term is the sum of all of the exponents of all of the variables. Don't forget that when we write a variable without an exponent, the exponent is 1. the degree of \(x^3y^2z = 3+2+1 = 6\)